a plane wavefront of 7000 A^0 Fallson an aperture. the area of half period zone of the diffraction pattern on screen 1 meter away from the aperture will be :
(1) 28*10^-7 m^2    (2) 44*10^-7 m^2    (3) 22*10^-7 m^2  (4) 14*10^-6

Grade:12

1 Answers

Arun
5 years ago
Dear student
Area = pi * p * lambda = 22/7 *1 * 7 * 10^-7 = 22 z 10^-7 m²
Regards
Arun (askIITians forum expert)